# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v2

@@ III.1 Models
  • $k$-local measurement operator $O^{k}$ used for measuring $k$-RDMs as given in Eq. (II), which are subsequently used to define PQKs according to Eq. (8)
@@
We point out that data encoding unitaries (cf. Eq. (3)) additionally have multiple degrees of freedom. Data encoding circuits are at the heart of each QML method and significantly influence core properties of the resulting model. Generating problem-specific encoding circuits with, e.g., proper gate sets and corresponding structure marks a distinct research branch, cf., e.g., Refs. [43, 24]. This is beyond the scope of this work, wherefore we restrict ourselves to nine data encoding circuits from the QML literature [44, 5, 7, 45, 31, 27, 28, 46]. In this context, we randomly initialize the variationally trainable parameters of those encoding circuits that involve $\bm{\theta}$ (cf. Eq. (3)) using a fixed seed. For details on these encoding circuits as well as corresponding illustrations, we refer to Appendix B.
III.2 Datasets
@@
two curves diff:
This dataset describes the curvature and distance of two one-dimensional curves embedded into a $d$-dimensional space. We follow the data generation procedure of Ref. [34] and use low-degree ($D$) Fourier series to embed two sets of data sampled from a 1-d interval as curves into $d$ dimensions, while adding Gaussian noise $\sigma=0.01$. The respective complexity can be controlled by fixing $d=4$ and vary $D=2,\ldots,20$, while adapting the offset $\Delta=\nicefrac{{1}}{{2D}}$.
hidden manifold diff:
This dataset is created by generating inputs on a low-dimensional manifold and label them by a simple neural network initialized at random. The inputs are then projected to a final $d$-dimensional space. We use the dataset generation procedure as described in Ref. [34] and vary the dimensionality $m$ of the manifold between $m=2,\ldots,20$ but keep the feature dimension constant at $d=4$.
Figure 3: Average feature correlations $\overline{C}$ of the $[0,1]$-normalized features to the outputs to assess the dataset complexity for the datasets considered in this study. The classification datasets depend on variables between $2$ and $20$ that can be seen as controlling the difficulty, while for regression datasets this variable corresponds to the number of features. Note that higher values of this measure indicate simpler problems.
The degree $D$ and the dimensionality $m$ of two curves diff and hidden manifold diff, respectively can be seen as control parameters adjusting the dataset complexity. To quantify this complexity, we use the average correlation $\overline{C}$ of the respective $[0,1]$-normalized features to the output as suggested in Ref. [49] and as implemented in the ECoL package [49, 50].
The results are displayed in Fig. 3. Note that the dataset difficulty is inverse to the $\overline{C}$ values, i.e., the lower means more difficult.
For regression, we select the following datasets:
Friedman #1:
The Friedman #1 regression problem is described in Refs. [51, 52]. The dataset consists of $d\geq 5$ independent features distributed uniformly on the interval $[0,1]$. The labels $y$ satisfy
QFMNIST:
This dataset is based on the fashion-MNIST dataset [53], which is send through a quantum circuit to create a quantum-based dataset. We use the procedure as described in Ref. [8]. For this, first the fashion-MNIST feature dimension is reduced by principle component analysis. The resulting $d$ features are then encoded into a quantum state using an $d$-qubit encoding circuit as proposed by Havlicek et al. [4]. The subsequent arbitrary single qubit rotations as applied in Ref. [8] are omitted. Finally, the labels are generated by computing expectation values w.r.t. measuring the first qubit in the Pauli-$Z$ basis. We generate datasets with $d=2,\ldots 15$ principal components.
@@
This real world dataset is taken from Ref. [54], which is a vast benchmark database for machine learning potential energy surfaces of a wide range of different molecules. Here, we use the ammonia data labeled with “STATIC-g32n1-1M” and “DZERO”. While the original dataset is in $xyz$-format we transformed them into internal coordinates (i.e. bond lengths and angles). Since nonlinear molecules with $N$ atoms show $f=3N-6$ degrees of freedom, this dataset consists of $d=6$ features with corresponding ground state energies (in Hartee) computed at CCSD(F12*)(T)/cc-pVDZ-F12 level of theory defining the labels. The dataset consists of 193 samples from which $M=155$ are used for training and $M^{\prime}=38$ for testing.
For the Friedman#1 and QFMNIST regression problems the number of features $d$ can be viewed as control parameter to adjust the dataset complexity (cf. Fig. 3).
III.3 Experimental setup and Implementation
@@
Design choices of quantum kernels
We use sQUlearn for evaluating quantum kernel Gram matrices, which provides FQKs according to Eq. (5) and allows for defining PQKs as generally given in Eq. (8). For PQKs, we investigate the impact of different outer kernel functions $\kappa$. Specifically, we consider the Gaussian (RBF) kernel (cf., Eq. (9), which is the default in sQUlearn), the Matérn kernel [41] with $\nu=\nicefrac{{3}}{{2}}$, i.e.,
$\kappa^{\text{Mat}}_{3/2}(\mathbf{x},\mathbf{x}^{\prime})$ $=\left(1+\frac{\sqrt{3}\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert}{\ell}\right)\exp\left(-\frac{\sqrt{3}\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert}{\ell}\right)\,,$ (12)
and the RationalQuadratic kernel [56], i.e.,
Moreover, we study the effect of using different $k$-local measurement operators $O^{k}$ for defining PQKs (cf., Eqs. (II) and (8)). We use the 1-RDM on all qubits w.r.t. different (combinations of) Pauli operators
$O^{k=1}$$\in\{X_{k=1},Z_{k=1},(X_{k=1}+Z_{k=1}),\sum_{P\in\mathcal{P}}P_{k=1}\}\,,$ (14)
@@
$P_{k=1}$ $=\sum_{k=1}^{n}P^{k}\,,$ (15)
where $P^{k}=\mathds{1}_{1}\otimes\dots\otimes P^{k}\otimes\mathds{1}_{k+1}\otimes\dots\otimes\mathds{1}_{n}$ is the Pauli operator $P$ acting on the $k$-th qubit. Additionally, we consider measuring the 2-RDM on all qubit combinations w.r.t. different Pauli operator configurations
$O^{k=2}$$\in\{X_{k=2},Z_{k=2},(X_{k=2}+Z_{k=2}),\sum_{P\in\mathcal{P}}P_{k=2}\}\,,$ (16)
where $P_{k=2}$ generally represents all possible permutations $S_{2}(n)$ of $2$-qubit Pauli measurements from $n$ qubits, i.e.
$P_{k=2}=\sum_{K\in S_{2}(n)}P^{K}\,,$ (17)
with $K$ the subset of $2$ qubits from $n$. Finally, we also check for one PQK definition with a combination of 1-RDM and 2-RDM measurements
@@
Hyperparameters of quantum kernel methods
The QKMs as implemented in sQUlearn work analogously to their classical counterparts in scikit-learn. In this work, we use QKRR and QSVR for regression tasks and QSVC for solving classification problems. The corresponding regularization hyperparameters are $\lambda$, $C$ and $\varepsilon$ as well as $C$, respectively as introduced above.
QKMTuner for hyperparameter search
We develop the tool QKMTuner to facilitate the extensive hyperparameter search of this study. The code is based on sQUlearn [31], Optuna [32] and scikit-learn [33]. QKMTuner consists of two main routines: a hyperparameter optimization within a grid search and a method for studying hyperparameter importances; cf. Fig. 1. The grid-search is given a list of data encoding circuits for each of which it builds a user-defined $n_{\mathrm{qubits}}\times n_{\mathrm{layers}}$-grid on which a hyperparameter search is performed for each grid point and each data encoding circuit for a given dataset. For evaluating hyperparameter importances, the respective method only takes a list of data encoding circuits and automatically searches for the optimal hyperparameters including $n_{\mathrm{qubits}}$ and $n_{\mathrm{layers}}$. In both cases, the hyperparameters are determined by maximizing the minimum between the mean and median of five-fold cross-validation scores. This choice of the objective function prevents the scores from becoming too optimistic and thus helps to prevent overfitting towards an easy fold. The respective scoring method can be set manually, whereas in this work we use the area under the receiver operating characteristic curve (ROC-AUC) from prediction scores for classification and the (negative) mean squared error (MSE) for regression tasks, respectively. For hyperparameter sampling we use the tree-structured parzen estimator [57], albeit different algorithms as provided by Optuna are supported. The code of QKMTuner is available via GitLab [30].
Data preprocessing
We scale the dataset features to the range $[f_{\text{min}},f_{\text{max}}]$ with $f_{\text{min}}\in[-\frac{\pi}{2},0)$ and $f_{\text{max}}\in(0,\frac{\pi}{2}]$ as in Ref. [34] and consider $f_{\text{max}}$ ($f_{\text{min}}$), or equivalently $w_{e}=f_{\text{max}}-f_{\text{min}}$, cf. Eq. (10), as additional hyperparameters which are optimized within QKMTuner [^3] This holds for all data encoding unitaries, except for the ChebyshevPQC, which encodes features non-linearly as $\arccos(x)$, wherefore we adapt $\tilde{f}_{\text{min}}\in[-1.0,0)$ and $\tilde{f}_{\text{max}}\in(0,1.0]$. For all regression datasets we additionally scale the corresponding target values to $[0,1]$. By using pipelines, we ensure that data scaling is part of the cross-validation step.
IV Results
In this section, we report all findings from the different types of experiments that have been performed, before we discuss these results in Sec. V.
IV.1 Model Performance
To start with, we delve into the relationships concerning the generalizability of QKMs. For studying this, we fix $n_{\mathrm{qubits}}$ for every dataset to the respective number of features for all upcoming investigations and perform hyperparameter optimizations for every data encoding circuit with $n_{\mathrm{layers}}\in[1,8]$
This analysis of test performances of different QKMs as a function of increasing dataset complexity is given in Fig. 4, which illustrates the results for the Friedman, QFMNIST, two curves diff, and hidden manifold diff datasets. In these cases, results for each QKM and dataset within the respective dataset family are aggregated across the data encoding circuits with corresponding obtained optimal $n_{\mathrm{layers}}^{*}$, yielding the best MSE or ROC-AUC score, respectively.
A comparison with Fig. 3 indicates that the behavior of test performance scores as a function of a control parameter qualitatively aligns with the increasing complexity of the respective datasets. This confirms that $\overline{C}$ is a useful measure of complexity. The comparison between FQK and PQK approaches in Figs. 4 (a) and (b), particularly for larger problem instances with increasing $d$, which transfers to increasing $n_{\mathrm{qubits}}$, reveals no significant performance differences. While PQKs demonstrate slightly superior performance, the difference is not as substantial as one might anticipate given the challenges posed in connection with exponential concentration [28], which might already be apparent as we approach 15 qubits. The absence of this observation may be explained by the presence of the bandwidth tuning parameter $w_{\mathrm{e}}$, which shows significant correlations with performance scores as discussed below. The comparison between QKRR and QSVR, each equipped with either a FQK or a PQK, does not reveal a clear advantage for either method. Notably, the QFMNIST ($d=2$) dataset constitutes the only case for which we observe a clear performance difference between FQK and PQK. Interestingly, this marks the most complex regression task, cf., Fig. 3. Whether this is merely a coincidence or actually due to a more fundamental mechanism might be interesting to investigate in the future. Regarding the classification tasks, we do not observe any significant difference between both QKMs.
We generally observe some outliers exhibiting particularly poor test performance scores. The comparison with associated training performance scores in Fig. A10 shows that this is probably a result of overfitting for certain model combinations. Since we investigate a large variety of combinations, it is unavoidable that some of the combinations result in models that are too expressive for some tasks. As shown in Appendix D.1 this is most dominant for QKRR-FQK/PQK approaches with increasing dataset complexity, while QSVR appears to be more robust. We point out here that our test size is quite small ($M^{\prime}=60$, as mentioned in Sec. III.2), which represents a clear limitation towards more detailed statements in this direction. This should be carefully considered in future studies.
Figure 4: Overview of test performance scores of respective QKMs as a function of increasing dataset complexity. Results within each dataset are aggregated across all data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$ yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays two regression tasks, where the MSE is used to measure the prediction accuracy. The Friedman dataset family is shown in (a), where the number of features can be viewed as control parameter for dataset complexity. The QFMNIST dataset family is shown in (b), where the number of principal components used in the dataset generation procedure controls the dataset complexity. The lower panel illustrates the two classification tasks of this study, where we use the ROC-AUC score to assess classification accuracy. In (c) we show the two curves diff dataset family with the degree $D$ controlling the complexity. The hidden manifold diff family is given in (d) with the manifold dimension $m$ as respective control parameter.
IV.2 Influence of Hyperparameters
@@
Figure 6: Comparison of hyperparameter importances for optimizing the five-fold cross-validation score in the corresponding hyperparameter searches for the classification tasks of this study. The results for each model are aggregated across all datasets within the correlating dataset family and all encoding circuits. The results corresponding to the two curves diff family are shown in (a). The hidden manifold diff dataset family is depicted in (b). In both cases we impose that $n_{\mathrm{qubits}}$ can only be integer multiples of the number of features ($d=4$), with a maximum of $n_{\text{qubits}}^{\text{max}}=15$.
@@
[…]
For classifications tasks, Fig. 6 displays the hyperparameter importances for the two curves diff and hidden manifold diff dataset families. Here, for both FQK and PQK QKMs, we aggregate results across all corresponding data subsets and encoding circuits. The results in Figs. 6 (a) and (b) clearly reveal that $w_{\mathrm{e}}$ is most important in order to obtain well trained quantum classifiers, which is in contrast to the regression results. Again, the importance of feature scaling can be directly related to the concept of kernel bandwidth tuning [26, 27] and is in accordance with the results presented therein. Moreover, it perfectly agrees with the recent outcomes of Ref. [37], where the scaling parameter $t$ of the Hamiltonian evolution feature map was evaluated as the most important parameter in PQKs for both the accuracy as well as for the geometric difference of selected classification tasks. In addition to that, we provide some evidence that the importance of feature prescaling is slightly more pronounced for FQKs than for PQKs. This, however, is not surprising, given the fact that the $\gamma$-parameter in PQKs of the form of Eq. (9) acts as an additional (classical) bandwidth tuning parameter. Here, $\gamma$ is the second most important parameter for PQK models. The regularization parameter $C$ of QSVC appears to be comparatively unimportant, except for some outliers. This is even more pronounced for PQKs than for FQKs. In comparison to Fig. 5, the results in Fig. 6 clearly indicate that $n_{\text{qubits}}$ and $n_{\text{layers}}$ appear to be more important, with tuning the number of layers being slightly more important than potentially encoding features redundantly on the number of qubits. This may be due to the increased dataset complexity for $D,m\geq 6$, cf. Fig. 3, which might require more expressive models. This can be achieved by introducing feature redundancies and adjusting the number of layers accordingly [11, 14].
@@
To obtain even deeper insights into the influence of hyperparameters, Fig. 7 provides an exemplary overview on the results of the correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves for the QFMNIST dataset family. We refer to Fig. A11 in Appendix D.2 for a summary of all other dataset families investigated in this study. The results are aggregated across datasets within a dataset family and across all nine encoding circuits and $n_{\mathrm{layers}}\in[1,8]$, respectively. As a measure for correlation, we calculate the Spearman coefficient, which neither requires normally distributed variables nor a linear relationship between them and is more robust to outliers. Figure 7 indicates statistically significant correlations by green-highlighted $p$-values. In line with the findings from the analysis of the importance of hyperparameter in Figs. 5 and 6, we observe moderate to strong statistically significant correlations between test performance scores and the various regularization parameters, i.e., $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively as well as $\varepsilon$ for QSVR. Additionally, consistent with findings on bandwidth tuning [26, 27], we frequently observe moderate and statistically significant correlations between the feature scaling parameter $w_{\mathrm{e}}$, cf. Eq. (10), and performance scores (also considering the findings from Appendix D.2 in Fig. A11). This parameter often also correlates with the $n_{\mathrm{layers}}$. For QSVR-/QSVC-PQK models, moderate correlations are also observed between the length-scale parameter $\gamma$ of the Gaussian outer kernel and test performance. Finally, we note that the $n_{\mathrm{layers}}$ parameter only shows weak statistically significant correlations to the test performance for the QFMNIST dataset.
@@
As mentioned above, the correlation analysis provides an explanation for the poor performance regimes observed in Fig. 4. We interpret the strong correlations between regularization parameters and model performance as indicative of ill-conditioned training kernel Gram matrices. This could be a result of unfavorable embeddings of the underlying data into the quantum Hilbert space. Furthermore, it corroborates the observed tendency to overfitting of some models.
IV.3 Influence of Encoding Circuits

### v3

@@ 3.1 Models
  • $k$-local measurement operator $O_{k}$ used for measuring $k$-RDMs as given in Eq. (2.2), which are subsequently used to define PQKs according to Eq. (7)
@@
We point out that data encoding unitaries (cf. Eq. (2)) additionally have multiple degrees of freedom. Data encoding circuits are at the heart of each QML method and significantly influence core properties of the resulting model. Generating problem-specific encoding circuits with, e.g., proper gate sets and corresponding structure marks a distinct research branch, cf., e.g., Refs. [2, 43]. This is beyond the scope of this work, wherefore we restrict ourselves to nine data encoding circuits from the QML literature [20, 42, 24, 30, 31, 11, 55, 26]. In this context, we randomly initialize the variationally trainable parameters of those encoding circuits that involve $\boldsymbol{\theta}$ (cf. Eq. (2)) using a fixed seed. For details on these encoding circuits as well as corresponding illustrations, we refer to Appendix C.
3.2 Datasets
@@
two curves diff This dataset describes the curvature and distance of two one-dimensional curves embedded into a $d$-dimensional space. We follow the data generation procedure of Ref. [8] and use low-degree ($D$) Fourier series to embed two sets of data sampled from a 1-d interval as curves into $d$ dimensions, while adding Gaussian noise $\sigma=0.01$. The respective complexity can be controlled by fixing $d=4$ and vary $D=2,\ldots,20$, while adapting the offset $\Delta=\nicefrac{{1}}{{2D}}$.
[…]
hidden manifold diff This dataset is created by generating inputs on a low-dimensional manifold and label them by a simple neural network initialized at random. The inputs are then projected to a final $d$-dimensional space. We use the dataset generation procedure as described in Ref. [8] and vary the dimensionality $m$ of the manifold between $m=2,\ldots,20$ but keep the feature dimension constant at $d=4$.
Figure 3: Average Spearman correlations $\overline{C}$ of all $[0,1]$-normalized features to the outputs to assess the dataset complexity for the datasets considered in this study. The classification datasets depend on variables between $2$ and $20$ that can be seen as controlling the difficulty, while for regression datasets this variable corresponds to the number of features. Note that higher values of this measure indicate simpler problems.
The degree $D$ and the dimensionality $m$ of two curves diff and hidden manifold diff, respectively can be seen as control parameters adjusting the dataset complexity. To quantify this complexity, we use the average of the (absolute value of the) Spearman correlations $\overline{C}$ of all respective $[0,1]$-normalized features to the output as suggested in Ref. [35] and as implemented in the ECoL package [35, 36].
The results are displayed in Fig. 3. Note that the dataset difficulty is inverse to the $\overline{C}$ values, i.e., the lower means more difficult.For these two classification tasks, we observe increasing complexity with decreasing qubit count.
[…]
QFMNIST This dataset is based on the fashion-MNIST dataset [59], which is send through a quantum circuit to create a quantum-based dataset. We use the procedure as described in Ref. [27]. For this, first the fashion-MNIST feature dimension is reduced by principle component analysis. The resulting $d$ features are then encoded into a quantum state using an $d$-qubit encoding circuit as proposed by Havlicek et al. [21]. The subsequent arbitrary single qubit rotations as applied in Ref. [27] are omitted. Finally, the labels are generated by computing expectation values w.r.t. measuring the first qubit in the Pauli-$Z$ basis. We generate datasets with $d=2,\ldots 15$ principal components.
[…]
NH3 Potential energy surface (PES) This real world dataset is taken from the bechmark database described in Ref. [45]. Here, we use the ammonia data labeled with “STATIC-g32n1-1M” and “DZERO” and transformed the $xyz$-data to internal coordinates (i.e. bond lengths and angles). Since nonlinear molecules with $N$ atoms show $f=3N-6$ degrees of freedom, this dataset consists of $d=6$ features with corresponding ground state energies (in Hartee). The dataset consists of 193 samples from which $M=155$ are used for training and $M^{\prime}=38$ for testing.
[…]
Within the hyperparameter optimization, we generally found that the computation time increases as the dataset complexity increases for both quantum and classical methods in classification and regression tasks. This is especially noticeable for (quantum) support vector machines for datasets with $\overline{C}<0.1$.
3.3 Experimental setup and Implementation
@@
[…]
Design choices of quantum kernels We use sQUlearn for evaluating quantum kernel Gram matrices, which provides FQKs according to Eq. (4) and allows for defining PQKs as generally given in Eq. (7). For PQKs, we investigate the impact of different outer kernel functions $\kappa$. Specifically, we consider the Gaussian (RBF) kernel (cf., Eq. (8), which is the default in sQUlearn), the Matérn kernel [44] with $\nu=\nicefrac{{3}}{{2}}$, cf. Eq. (A.1), and the RationalQuadratic kernel [15], cf. Eq. (19)
Moreover, we study the effect of using different $k$-local measurement operators $O_{k=w(O)}$ [^3] Here we denote $w(O)$ the weight of the observable $O$, i.e., the number of qubits on which it acts nontrivially. for defining PQKs (cf., Eqs. (2.2) and (7)). We use the 1-RDM on all qubits w.r.t. different (combinations of) Pauli operators
$O_{k=1}$$\in\{X_{k=1},Z_{k=1},(X_{k=1}+Z_{k=1}),\sum_{P\in\mathcal{P}}P_{k=1}\}\,,$ (10)
@@
$P_{k=1}$ $=\sum_{i=1}^{n}P_{i}\,,$ (11)
where $P_{i}=\mathds{1}_{1}\otimes\dots\otimes P_{i}\otimes\mathds{1}_{i+1}\otimes\dots\otimes\mathds{1}_{n}$ is the Pauli operator $P$ acting on the $i$-th qubit. Additionally, we consider measuring the 2-RDM on all qubit combinations w.r.t. different Pauli operator configurations
$O_{k=2}$$\in\{X_{k=2},Z_{k=2},(X_{k=2}+Z_{k=2}),\sum_{P\in\mathcal{P}}P_{k=2}\}\,,$ (12)
where $P_{k=2}$ generally represents all possible permutations of $2$-qubit Pauli measurements from $n$ qubits, i.e.
$P_{k=2}=\sum_{i<j}P_{i}P_{j}\,.$ (13)
Finally, we also check for one PQK definition with a combination of 1-RDM and 2-RDM measurements
@@
Hyperparameters of quantum kernel methods The QKMs as implemented in sQUlearn work analogously to their classical counterparts in scikit-learn. In this work, we use QKRR and QSVR for regression tasks and QSVC for solving classification problems. The corresponding regularization hyperparameters are $\lambda$, $C$ and $\varepsilon$ as well as $C$, respectively as introduced above. While QKRR and QSVR leverage the same quantum kernel, they differ in the formulation of their classical kernel method; cf. Appendix A.2. Considering both is essential to choose the best approach for a given dataset characteristic.
[…]
QKMTuner for hyperparameter search We develop the tool QKMTuner to facilitate the extensive hyperparameter search of this study. The code is based on sQUlearn [31], Optuna [1] and scikit-learn [41]. QKMTuner consists of two main routines: a hyperparameter optimization within a grid search and a method for studying hyperparameter importances; cf. Fig. 2. Details on the implementation and how to use QKMTuner are outlined in Appendix E. The code is available via GitLab [46], where we provide additional documentation.
[…]
Data preprocessing We scale the dataset features to the range $[f_{\text{min}},f_{\text{max}}]$ with $f_{\text{min}}\in[-\frac{\pi}{2},0)$ and $f_{\text{max}}\in(0,\frac{\pi}{2}]$ as in Ref. [8] and consider $f_{\text{max}}$ ($f_{\text{min}}$), or equivalently $w_{e}=f_{\text{max}}-f_{\text{min}}$, cf. Eq. (9), as additional hyperparameters which are optimized within QKMTuner [^4] This holds for all data encoding unitaries, except for the ChebyshevPQC, which encodes features non-linearly as $\arccos(x)$, wherefore we adapt $\tilde{f}_{\text{min}}\in[-1.0,0)$ and $\tilde{f}_{\text{max}}\in(0,1.0]$. For all regression datasets we additionally scale the corresponding target values to $[0,1]$. By using pipelines, we ensure that data scaling is part of the cross-validation step.
4 Results
In this section, we report all findings from the different types of experiments that have been performed, before we discuss these results in Sec. 5.
4.1 Model Performance
Figure 4: Overview of test performance scores of respective QKMs as a function of increasing dataset complexity. Results within each dataset are aggregated across all data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$ yielding minimum/maximum test performance scores for regression/classification, respectively. For comparison, we provide classical KRR/SVR and SVC results each based on a RBF kernel. The upper panel displays two regression tasks, where the MSE is used to measure the prediction accuracy and the number of features controls the dataset complexity. The Friedman dataset family is shown in (a). The QFMNIST dataset family is shown in (b). The lower panel illustrates the two classification tasks of this study, where we use the ROC-AUC score to assess classification accuracy. In (c) we show the two curves diff dataset family with the degree $D$ controlling the complexity. The hidden manifold diff family is given in (d) with the manifold dimension $m$ as respective control parameter.
To start with, we delve into the relationships concerning the generalizability of QKMs. For studying this, we fix $n_{\mathrm{qubits}}$ for every dataset to the respective number of features for all upcoming investigations and perform hyperparameter optimizations for every data encoding circuit with $n_{\mathrm{layers}}\in[1,8]$ To gauge the performance of QKMs, we also provide classical results with KRR/SVR and SVC based on RBF kernels. The corresponding hyperameters were optimized with Optuna using the same parameter ranges as introduced in Sec. 3.3 for QKMs.
This analysis of test performances as a function of increasing dataset complexity is given in Fig. 4, which illustrates the results for the Friedman, QFMNIST, two curves diff, and hidden manifold diff datasets. In these cases, results for each QKM and dataset within the respective dataset family are aggregated across the data encoding circuits with corresponding obtained optimal $n_{\mathrm{layers}}^{*}$, yielding the best MSE or ROC-AUC score, respectively.
A comparison with Fig. 3 indicates that the behavior of test performance scores qualitatively aligns with the increasing complexity of the respective datasets. This confirms that $\overline{C}$ is a useful measure of complexity. The comparison between FQK and PQK approaches in Figs. 4 (a) and (b), particularly for larger problem instances with increasing $d$, which transfers to increasing $n_{\mathrm{qubits}}$, reveals no significant performance differences. While PQKs demonstrate slightly superior performance, the difference is not as substantial as one might anticipate given the challenges posed in connection with exponential concentration [55], which might already be apparent as we approach 15 qubits. The absence of this observation may be explained by the presence of the bandwidth tuning parameter $w_{\mathrm{e}}$, which shows significant correlations with performance scores as discussed below. The comparison between QKRR and QSVR, each equipped with either a FQK or a PQK, does not reveal a clear advantage for either method. Notably, the QFMNIST ($d=2$) dataset constitutes the only case for which we observe a clear performance difference between FQK and PQK. Interestingly, this marks the most complex regression task, cf., Fig. 3. Whether this is merely a coincidence or actually due to a more fundamental mechanism might be interesting to investigate in the future. Regarding the classification tasks, we do not observe any significant difference between both QKMs.
We generally observe some outliers exhibiting particularly poor test performance scores. The comparison with associated training performance scores in Fig. A10 shows that this is probably a result of overfitting for certain model combinations. Since we investigate a large variety of combinations, it is unavoidable that some of the combinations result in models that are too expressive for some tasks. As shown in Appendix G.1 this is most dominant for QKRR-FQK/PQK approaches with increasing dataset complexity, while QSVR appears to be more robust.
The classical baselines follow the general trend of QKMs for the regression tasks. In the classification examples, the RBF-SVC results align with the overall difficulty trends of both problems and face challenges with the same datasets as the quantum models. In some instances, classical models outperform QKMs, but they perform poorly in the hidden manifold diff family, with ROC-AUC values below 0.5. This is due to significant overfitting in SVC models, as noted in Ref. [8], with overfitting sometimes more severe in classical baselines. This may be due to dataset size limitations or using default classical models. Future studies should explore classical simulability [53] and the impact of providing more data [22], to assess requirements for quantum advantage and associated geometric differences.
4.2 Influence of Hyperparameters
[…]
@@ 4.2 Influence of Hyperparameters
Figure 6: Comparison of hyperparameter importances for optimizing the five-fold cross-validation score in the corresponding hyperparameter searches for the classification tasks of this study. The results for each model are aggregated across all datasets within the correlating dataset family and all encoding circuits. We chose half the interquartile range to define the whisker length. The results corresponding to the two curves diff family are shown in (a). The hidden manifold diff dataset family is depicted in (b). In both cases we impose that $n_{\mathrm{qubits}}$ can only be integer multiples of the number of features ($d=4$), with a maximum of $n_{\text{qubits}}^{\text{max}}=15$.
@@
For classifications tasks, Fig. 6 displays the hyperparameter importances. Here, for both FQK and PQK QKMs, we aggregate results across all corresponding data subsets and encoding circuits. The results clearly reveal that $w_{\mathrm{e}}$ is most important in order to obtain well trained quantum classifiers, which is in contrast to the regression results. Again, this can be directly related to the concept of kernel bandwidth tuning and is in accordance with the results presented therein. Moreover, it perfectly agrees with the recent outcomes of Ref. [16]. In addition to that, we provide some evidence that the importance of feature prescaling is slightly more pronounced for FQKs than for PQKs. This, however, is not surprising, given the fact that the $\gamma$-parameter in PQKs of the form of Eq. (8) acts as an additional (classical) bandwidth tuning parameter. Here, $\gamma$ is the second most important parameter for PQK models. The regularization parameter $C$ of QSVC appears to be comparatively unimportant, except for some outliers. This is even more pronounced for PQKs than for FQKs. In comparison to Fig. 5, the results in Fig. 6 clearly indicate that $n_{\text{qubits}}$ and $n_{\text{layers}}$ appear to be more important, with tuning the number of layers being slightly more important than potentially encoding features redundantly on the number of qubits. This may be due to the increased dataset complexity for $D,m\geq 6$, cf. Fig. 3, which might require more expressive models. This can be achieved by introducing feature redundancies and adjusting the number of layers accordingly [49, 51].
@@
To obtain even deeper insights into the influence of hyperparameters, Fig. 7 provides an exemplary overview on the results of the correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves for the QFMNIST dataset family. We refer to Fig. A11 in Appendix G.2 for a summary of all other dataset families investigated in this study. The results are aggregated across datasets within a dataset family and across all nine encoding circuits and $n_{\mathrm{layers}}\in[1,8]$, respectively. We calculate the Spearman coefficient to measure correlaiton. Figure 7 indicates statistically significant correlations by green-highlighted $p$-values. In line with the aforementioned findings from the analysis of the importance of hyperparameter, we observe moderate to strong statistically significant correlations between test performance scores and the various regularization parameters, i.e., $\lambda$ and $C$ for QKRR and QSVR/QSVC, respectively as well as $\varepsilon$ for QSVR. Additionally, consistent with findings on bandwidth tuning [52, 11], we frequently observe moderate and statistically significant correlations between the feature scaling parameter $w_{\mathrm{e}}$, cf. Eq. (9), and performance scores (also considering the findings from Appendix G.2 in Fig. A11). This parameter often also correlates with the $n_{\mathrm{layers}}$. For QSVR-/QSVC-PQK models, moderate correlations are also observed between the length-scale parameter $\gamma$ of the Gaussian outer kernel and test performance. Finally, we note that the $n_{\mathrm{layers}}$ parameter only shows weak statistically significant correlations to the test performance for the QFMNIST dataset.
@@
As mentioned above, the correlation analysis provides an explanation for the poor performance regimes observed in Fig. 4. We interpret the strong correlations between regularization parameters and model performance as indicative of ill-conditioned training kernel Gram matrices. Furthermore, it corroborates the observed tendency to overfitting of some models.
4.3 Influence of Encoding Circuits